A trapezoid has an area of 36 sq. m and an altitude of 2m. Its two bases have ratio of 4:5. Find the perimeter of the trapezoid. Find the percentage increase in the volume if its edge is increased by 1%


1
Expert's answer
2022-03-19T02:40:19-0400

"S=36m^2\\
BK=2m\\
BC:AD=4:5"

BK, CM - an altitude

Lets

"BC=4x, AD=5x\\
S=\frac{BC+AD}{2}\cdot BK\\
S=\frac{4x+5x}{2}\cdot2=36\\
9x=36\\
x=4\\
BC=16m, AD=20m"

"AK=MD=\frac{AD-BC}{2}=\frac{20-16}{2}=2(m)\\"

"\triangle ABK, \angle K=90^0\\
AB^2=AK^2+BK^2\\
AB^2=2^2+2^2=4+4=8\\
AB=\sqrt{8}=2\sqrt{2}(m)\\
AB=CD=2\sqrt{2}m\\
P=AB+BC+CD+AD=\\
=2\sqrt{2}+16+2\sqrt{2}+20=36+4\sqrt{2}(m)"

If AB and CD is increased by 1%=0.01:

"AB'=1.01(2\sqrt{2})=2.02\sqrt{2}(m)\\
P'=16+20+2.02\sqrt{2}+2.02\sqrt{2}=36+4.04\sqrt{2}(m)\\
\frac{P'\cdot100\%}{P}=\frac{(36+4.04\sqrt{2})\cdot100\%}{36+4\sqrt{2}} \approx100.14\%"